# Transfer Learning for Spatial Autoregressive Models

### v1

@@ 4 The debiasing stage:
In particular, we consider a classical 2SLS procedure in the second stage as an example, which has been widely used for the SAR model to deal with the endogeneity. Given an $n\times d$ ($d\geq p+q$) instrumental variables matrix $Q^{(0)}$ containing $\mathbf{X}^{(0)}$, the first-stage estimator in the 2SLS procedure is
$\widehat{\pi}_{j}=\underset{\pi_{j}\in\mathbb{R}^{d}}{\arg\min}\left\{\frac{1}{2n_{0}}\|{\mathbb{X}}^{(0)}_{j}-Q^{(0)}\pi_{j}\|_{2}^{2}\right\},$ (2.7)
for $j\in[p+q]$, where ${\mathbb{X}}^{(0)}=({W_{1}}^{(0)}Y^{(0)},\dots,{W_{p}}^{(0)}Y^{(0)},\mathbf{X}^{(0)})\in\mathbb{R}^{p+q}$, ${\mathbb{X}}^{(0)}_{j}$ is the $j$-th column of ${\mathbb{X}}^{(0)}$ and $\pi_{j}$’s are coefficients of instrumental variables. Denote the fitted values of ${\mathbb{X}}^{(0)}_{j}$’s by $\widehat{\mathbb{X}}^{(0)}_{j}\equiv Q^{(0)}\hat{\pi}_{j}=P_{Q^{(0)}}{\mathbb{X}}^{(0)}_{j}$, and $\widehat{\mathbb{X}}^{(0)}=(\widehat{\mathbb{X}}^{(0)}_{1},\dots,\widehat{\mathbb{X}}^{(0)}_{p+q})$, where the projection matrix is $P_{Q^{(0)}}\equiv{Q^{(0)}}({Q^{(0)}}^{\intercal}Q^{(0)})^{-1}{Q^{(0)}}^{\intercal}$. In the SAR model, instrumental variables could be $Q^{(0)}=(\mathbf{X}^{(0)},W_{1}^{(0)}\mathbf{X}^{(0)},\dots,W_{p}^{(0)}\mathbf{X}^{(0)})$ with $d=q(1+p)$ columns. We note that if the first column of $\mathbf{X}^{(0)}$ is the intercept term 1, we need to drop it in $W_{j}^{(0)}\mathbf{X}^{(0)}(j\in[p])$ to avoid multicollinearity. See Kelejian and Prucha, (1998) and Lee, (2003) for several alternative choices of instruments. The corresponding loss function for the second stage estimation is
$\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\theta\right)=\|Y^{(0)}-\widehat{\mathbb{X}}^{(0)}\theta\|_{2}^{2}.$ (2.8)
@@
$\widehat{\theta}_{\mathcal{A}\textit{-TranSAR}}=\underset{\theta\in\mathbb{R}^{p+q}}{\arg\min}\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\theta\right)+\lambda_{\delta}\|\theta-\widehat{\omega}\|_{1}.$ (2.9)
@@
$\widehat{\pi}_{j}=\underset{\pi_{j}\in\mathbb{R}^{d}}{\arg\min}\left\{\frac{1}{2n_{0}}\|{\mathbb{X}}^{(0)}_{j}-Q^{(0)}\pi_{j}\|_{2}^{2}+p_{\lambda_{\pi_{j}}}(\pi_{j})\right\},$ (2.10)
@@ 2.3 A transferable source detection algorithm
An intuitive approach might involve seeking a metric of similarity between each source and the target data. We compute the initial loss based on the target dataset. Subsequently, we combine the target dataset with each source dataset individually to estimate and compute the corresponding losses. If a source dataset is similar to the target dataset, we anticipate two loss values to be comparably close. Hence, we select a critical value to identify which source datasets are informative. For spatial datasets characterized by a network structure, data splitting used in Tian and Feng, (2023) can not be directly applied since it could disrupt this network integrity. As a result, we adopt the spatial residual bootstrap approach as an alternative. This idea is illustrated in details as follows.
@@
$Y^{(0,r)}=$ $\mathbb{X}^{(0)}\widehat{\theta}_{ini}+\widehat{e}^{(r)},~{}~{}\mathbf{X}^{(0,r)}=$ $\mathbf{X}^{(0)},$ (2.11)
for $r=1,2,3$. Let $Z^{(0,r)}=(Y^{(0,r)},\mathbf{X}^{(0,r)})$ for $r=1,2,3$. Let $Z^{(0,-r)}$ be the concatenated bootstrap samples without the $r$the sample, for example, $Z^{(0,-1)}={({Z^{(0,2)}}^{\intercal},{Z^{(0,3)}}^{\intercal})}^{\intercal}$. Define $\mathbf{X}^{(0,-r)},\mathbb{X}^{(0,r)}$ and $\mathbb{X}^{(0,-r)}$ similarly.
Second, for each $r=1,2,3$, we fit the model in (2.1) on the catenated bootstrap samples $Z^{(0,-r)}$ and use the resulting estimator $\widehat{\theta}^{(0,r)}$ to compute the loss value on the remaining bootstrap sample $Z^{(0,r)}$, that is
$\mathcal{L}^{[r]}(\widehat{\theta}^{(0,r)})\equiv\frac{1}{2n_{0}}\|Y^{(0,r)}-\mathbb{X}^{(0,r)}\widehat{\theta}^{(0,r)}\|_{2}^{2}.$ (2.12)
@@
$\hat{\sigma}=\sqrt{\frac{1}{2}\sum_{r=1}^{3}{(\mathcal{L}^{[r]}(\hat{\theta}^{(0,r)})-\mathcal{L}^{(0)})}^{2}},$ (2.13)
which measures the dispersion of bootstrapped loss values $\mathcal{L}^{[r]}(\widehat{\theta}^{(0,r)})$ relative to their average $\mathcal{L}^{(0)}$. Next, for each $r=1,2,3$, we combine $Z^{(0,-r)}$ with each $k$-th source dataset, fit the model in (2.1) on the newly combined data and compute the loss denoted by $\mathcal{L}^{(k,r)}$ on $Z^{(0,r)}$. Similarly, the average loss for the $k$-th source dataset over three-fold bootstrap samples is $\mathcal{L}^{(k)}$. Consequently, the difference between $\mathcal{L}^{(k)}$ and $\mathcal{L}^{(0)}$ provides a metric of the similarity between the $k$-th source dataset and the target data. Source datasets whose corresponding loss differences below certain threshold are recognized as transferable informative data. Following the idea of Tian and Feng, (2023), we choose $\hat{\sigma}\vee 0.01$ as the detection threshold. Then, the estimated transferable source index set $\widehat{\mathcal{A}}$ is defined as
$\widehat{\mathcal{A}}=\left\{k=1,\ldots,K\mid\left|\mathcal{L}^{(k)}-\mathcal{L}^{(0)}\right|\leq\left(\widehat{\sigma}\vee 0.01\right)\right\}.$ (2.14)
@@ 4 Residuals: $\hat{e}=Y^{(0)}-\mathbb{X}^{(0)}\hat{\theta}_{ini}$ ;
21The estimated transferable source index set: $\hat{\mathcal{A}}\leftarrow\left\{k=1,\ldots,K\mid\left|\mathcal{L}^{(k)}-\mathcal{L}^{(0)}\right|\leq\left(\hat{\sigma}\vee 0.01\right)\right\}$;
@@
3 Theoretical results
In Section 3, we will establish an estimation convergence rate for the proposed $\mathcal{A}$-TranSAR estimator by assuming $\underline{n}\equiv\min_{k\in[K]\cup\{0\}}n_{k}\rightarrow\infty$. Subsection 3.1 is for general results and Subsection 3.2 is based on the 2SLS estimation method. Finally, we prove the consistency of the transferable source detection algorithm in Subsection 3.3.
For theoretical analysis, we define the population form for the above estimator is defined as
$\omega^{\mathcal{A}}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\sum_{k\in\mathcal{A}}\alpha_{k}\bar{\mathit{\ell}}_{1}^{(k)}(\omega),$ (3.1)
[…]
$\omega^{(k)}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\ \bar{\mathit{\ell}}_{1}^{(k)}(\omega).$ (3.2)
3.1 Convergence rate when $\mathcal{A}_{h}$ is known
Assumption 1 (Network structure).
The spatial weight matrices $W_{l}^{(k)}$, for $k=0,1,\ldots,K,l=1,\dots,p$, are non-stochastic with zero diagonals, and $p$ is fixed and does not depend on $n_{0}$ or $n_{k}$. $S^{(k)}(\lambda)=I_{n_{k}}-\sum_{l=1}^{p}\lambda_{l}^{(k)}W_{l}^{(k)}$ is invertible for all $\lambda\in\Lambda$, where the parameter space $\Lambda$ is compact. Matrices $W_{l}^{(k)}$ and $S^{(k)}(\lambda)^{-1}$ are uniformly bounded in both row and column sum norms for all $n_{k},k\in[K]\cup\{0\},l\in[p]$ and $\lambda\in\Lambda$. And for all large enough $n_{k}$, $\|W_{l}^{(k)}S^{(k)}(\lambda)^{-1}\|_{F}/\sqrt{n_{k}}$ is bounded away from zero and infinity uniformly over $k\in[K]\cup\{0\},l\in[p]$ and $\lambda\in\Lambda$.
(i) Denote $\mathbf{X}^{(0)}\equiv{(\mathbf{x}_{1}^{(0)},\dots\mathbf{x}_{n_{0}}^{(0)})}^{\intercal}$ and the covariance matrix of $\mathbf{x}_{i}^{(0)}$ for each $i$ is $\Sigma_{0}$. For any ${\beta}\in\mathbb{R}^{q},{\mathbf{x}_{i}^{(0)}}^{\intercal}{\beta}\ (i\in[n_{0}])$, are $i.i.d.$ $\kappa_{0}\|{\beta}\|_{2}$ sub-Gaussian with mean zero, and $0<\underline{\psi_{0}}=\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})=\overline{\psi_{0}}<\infty$. (ii) The error terms $V_{i}^{(0)}\ (i\in[n_{0}])$ are $i.i.d.$sub-Gaussian with mean zero and variance $\sigma^{2}$. (iii) As $n_{0}$rightarrow +$\infty$+ ∞, $\mathbb{P}\left(\lambda_{\min}(\frac{(\widehat{\mathbb{X}}^{(0)})^{\intercal}\widehat{\mathbb{X}}^{(0)}}{n_{0}})\geq\phi_{0}\right)\rightarrow 1$, where $\phi_{0}$ is a positive constant.
Assumption 3 (Convergence rates of the first step estimator).
The first step estimator $\widehat{\omega}$ minimizing (2.4) satisfies $\|\widehat{\omega}-\omega^{\mathcal{A}}\|_{1}=O_{\mathbb{P}}{(a_{n_{\mathcal{A}}}^{(1)})}$, and $n_{\mathcal{A}}^{-1}\|{{\mathbb{X}}^{\mathcal{A}}}(\widehat{\omega}-\omega^{\mathcal{A}})\|_{2}^{2}=O_{\mathbb{P}}(a_{n_{\mathcal{A}}}^{(2)})$, where $a_{n_{\mathcal{A}}}^{(1)}\rightarrow 0$ and $a_{n_{\mathcal{A}}}^{(2)}\rightarrow 0$, as $n_{\mathcal{A}}\rightarrow\infty$, and ${\mathbb{X}}^{\mathcal{A}}$ is stack of $\mathbb{X}^{(k)}$ for $k\in\mathcal{A}$.
Assumption 4 (Instruments).
When $n_{0}$ is large enough, the IV matrix $Q^{(0)}=\left(\mathbf{X}^{(0)},W_{1}^{(0)}\mathbf{X}^{(0)},\ldots,W_{q}^{(0)}\mathbf{X}^{(0)}\right)$ satisfies $0<\psi_{*}^{(0)}\leq\frac{1}{\sqrt{n_{0}}}\|Q^{(0)}\|_{2}\leq{\psi^{(0)}}^{*}<\infty$ for some constants $\psi_{*}^{(0)}$ and ${\psi^{(0)}}^{*}$. There exist constant coefficient vectors ${\pi}_{j}\in\mathbb{R}^{d}$ for $j\in[p+q]$ such that $\max_{j=1,\dots,p+q}\|\widehat{\pi}_{j}-{\pi}_{j}\|_{2}=O_{\mathbb{P}}(\sqrt{\log(d)/n_{0}})$ and $\|\pi_{j}\|_{2}^{2}$ is uniformly bounded over $j$ and $n_{0}$, where $\widehat{\pi}_{j}\ (j\in[p+q])$ are defined in Eq. (2.7).
Remark 4.
Theorem 1 (Convergence rate of the $\mathcal{A}$-TranSAR estimator).
Let the parameter space be $\Theta=\Lambda\otimes\left\{{\beta}\mid\|{\beta}\|_{2}<M\right\}$ for some positive constant $M$. Under Assumptions 1-4, if the loss function $\operatorname{\mathit{\ell}}_{1}$ is twice differentiable, $\sup_{k\in\mathcal{A}}\|{\Sigma^{\mathcal{A}}}^{-1}\Sigma^{(k)}\|_{1}=\check{C}<\infty$, $s\frac{\log{q}}{n_{0}}=o(1)$, $q\asymp d$, and $\lambda_{\delta}\gtrsim\sqrt{\log{(q)}/n_{0}}$, where
[…]
Denote $S=\operatorname{support}(\theta_{0})$ and $s=|S|$. Then
$\|\widehat{\theta}-\theta_{0}\|_{2}^{2}=O_{\mathbb{P}}\left((\sqrt{\frac{\log{q}}{n_{0}}}h)\wedge(s\frac{\log{q}}{n_{0}})\wedge h^{2}+a_{n_{\mathcal{A}}}^{(2)}\right).$ (3.5)
Remark 5.
3.2 Convergence rate under the 2SLS loss
In this section, we verify the requirements in the above theorem for the 2SLS estimation and simplify the convergence rate of the $\mathcal{A}$-TranSAR estimator. As in Eq. (2.8), let $\operatorname{\mathit{\ell}}_{1}(\cdot\mid\omega)$ be the loss function of the 2SLS. To elaborate, let $Y^{\mathcal{A}}$ denote the stack consisting of $Y^{(k)}$ for all $k\in\mathcal{A}$, i.e., $Y^{\mathcal{A}}={({Y^{(k)}}^{\intercal},k\in\mathcal{A})}^{\intercal}$. $\mathbf{X}^{\mathcal{A}}$ and ${\mathbb{X}}^{\mathcal{A}}$ are defined similarly. Then the objective function (2.4) could be rewritten as
$\frac{1}{2n_{\mathcal{A}}}\|Y^{\mathcal{A}}-\widehat{\mathbb{X}}^{\mathcal{A}}\omega\|_{2}^{2}+\lambda_{\omega}\|\omega\|_{1},$ (3.6)
where $\widehat{\mathbb{X}}^{(k)}$ is defined similarly as $\widehat{\mathbb{X}}^{(0)}$ in Subsection 2.2, and $\widehat{\mathbb{X}}^{\mathcal{A}}$ is their stack. Here the IV matrices are $Q^{(k)}$ and $Q^{\mathcal{A}}$. We impose some several similar conditions on $\mathfrak{D}^{(k)}$, $k\in\mathcal{A}$ in the following.
Assumption 5 (Regular conditions for design matrices and error terms on the $k$-th source).
For source $k\in\mathcal{A}$, we have the following assumptions: (i) For the design matrix $\mathbf{X}^{(k)}=(\mathbf{x}_{1}^{(k)},\dots\mathbf{x}_{n_{k}}^{(k)})^{\intercal}$ and any ${\beta}\in\mathbb{R}^{q},{\mathbf{x}_{i}^{(k)}}^{\intercal}{\beta}\ (i\in[n_{k}])$ are $i.i.d.$ $\kappa_{k}\|{\beta}\|_{2}$ sub-Gaussian with mean zero for certain constant $\kappa_{k}$, where $\max_{k\in\mathcal{A}}\kappa_{k}<\infty$. (ii) The error terms $V_{i}^{(k)}\ (i\in[n_{k}])$ are $i.i.d.$sub-Gaussian with mean zero and variance $\sigma_{k}^{2}$. (iii) As $n_{k}\rightarrow+\infty$, $\mathbb{P}\left(\lambda_{\min}\left(\frac{{{\widehat{\mathbb{X}}}^{(k)\intercal}}{\widehat{\mathbb{X}}}^{(k)}}{n_{k}}\right)\geq\phi_{k}\right)\rightarrow 1$, where $\phi_{k}$ is a positive constant.
Assumption 6 (Instruments for the $k$-th source).
For the instruments
$Q^{(k)}=\left(\mathbf{X}^{(k)},W_{1}^{(k)}\mathbf{X}^{(k)},\ldots,W_{q}^{(k)}\mathbf{X}^{(k)}\right),\ k\in\mathcal{A}\cup\{0\},$ (3.7)
when $n_{k}$, the number of rows of $Q^{(k)}$, is large enough, $0<\psi_{*}^{(k)}\leq\frac{1}{\sqrt{n_{k}}}\|Q^{(k)}\|_{2}\leq{\psi^{(k)}}^{*}<\infty$ for some constants $\psi_{*}^{(k)}$ and ${\psi^{(k)}}^{*}$. There exist constant coefficient vectors $\pi_{j}^{(k)}\in\mathbb{R}^{d}$ for $j\in[p+q]$ such that $\max_{j=1,\dots,p+q}\|\hat{\pi}_{j}-\pi_{j}^{(k)}\|_{2}=O_{\mathbb{P}}(\sqrt{\log(d)/n_{k}})$ and $\|\pi_{j}^{(k)}\|_{2}^{2}$ is uniformly bounded over $j$ and $k$.
Theorem 2.
$\|\widehat{\theta}-\theta_{0}\|_{2}^{2}=O_{\mathbb{P}}\left((\sqrt{\frac{\log{q}}{n_{0}}}h)\wedge({\frac{s\log{q}}{n_{0}}})\wedge h^{2}+(\frac{\log{n_{A}}}{n_{\mathcal{A}}}h^{2})\wedge\frac{\log^{3}{n_{A}}}{n_{\mathcal{A}}^{3}}+\frac{s\log{q}}{n_{\mathcal{A}}}\right).$ (3.8)
3.3 Detection consistency
Assumption 7 (Identifiability condition).
The tolerance level $h=o(1)$ as $\underline{n}\equiv\min_{k\in[K]\cup\{0\}}n_{k}\rightarrow\infty$. For any $k\in[K]\cup\{0\}$ and $r\in[3]$, $q\log(q)\|\widetilde{\theta}^{(k)}-\widehat{\theta}^{(k,r)}\|_{1}=o_{\mathbb{P}}(1)$. For $k\in\mathcal{A}^{c}\equiv[K]-\mathcal{A}$, we control the signal power as
$\frac{\sigma^{2}}{2n_{0}}\operatorname{tr}\{[(S(\tilde{\lambda}^{(k)})-S)S^{-1}]^{\intercal}[(S(\tilde{\lambda}^{(k)})+S)S^{-1}]\},$ (3.9)
and the gap between $\theta_{0}$ and $\widetilde{\theta}^{(k)}$,
$\varphi_{k}\equiv\sqrt{\frac{1}{n_{0}}}(\|\theta_{0}-\widetilde{\theta}^{(k)}\|_{1}+1)^{2}=o(1),$ (3.10)
Remark 6.
Conditions (7) and (3.10) guarantee a suitable gap between the population-level coefficients in set $\mathcal{A}^{c}$ and the true coefficient of the target model. Intuitively, assuming an appropriate signal strength is necessary to ensure the consistency of estimators. On the one hand, a very weak signal makes it challenging for the algorithm to differentiate informative and uninformative sources. On the other hand, an excessively strong signal can lead the first step estimator to mimic the model structure of the source, thereby hindering accurate identification.
Theorem 3 (Detection consistency of $\widehat{\mathcal{A}}$).
If Assumptions (1)-(4) and (7) are satisfied, $\|\beta_{0}\|_{2}<\infty$ and $\frac{\log q}{n_{0}}=o(1)$, then the automatically detected transferable index set $\widehat{\mathcal{A}}$ in Eq. (2.14) is consistent, i.e.,
[…]

### v2

@@ 4 The debiasing stage:
In particular, we consider a classical 2SLS procedure in the second stage as an example, which has been widely used for the SAR model to deal with the endogeneity. Given an $n_{0}\times d$ ($d\geq p+q$) instrumental variables matrix $Q^{(0)}$ containing $\mathbf{X}^{(0)}$, the first-stage estimator in the 2SLS procedure is
$\widehat{\pi}_{j}=\underset{\pi_{j}\in\mathbb{R}^{d}}{\arg\min}\left\{\frac{1}{2n_{0}}\|{\mathbb{X}}^{(0)}_{j}-Q^{(0)}\pi_{j}\|_{2}^{2}\right\},$ (7)
for each $j\in[p+q]$, where ${\mathbb{X}}^{(0)}=({W_{1}}^{(0)}Y^{(0)},\dots,{W_{p}}^{(0)}Y^{(0)},\mathbf{X}^{(0)})\in\mathbb{R}^{n_{0}\times(p+q)}$, ${\mathbb{X}}^{(0)}_{j}$ is the $j$-th column of ${\mathbb{X}}^{(0)}$ and $\pi_{j}$’s are coefficients of instrumental variables. Denote the fitted values of ${\mathbb{X}}^{(0)}_{j}$’s by $\widehat{\mathbb{X}}^{(0)}_{j}\equiv Q^{(0)}\hat{\pi}_{j}=P_{Q^{(0)}}{\mathbb{X}}^{(0)}_{j}$, and $\widehat{\mathbb{X}}^{(0)}=(\widehat{\mathbb{X}}^{(0)}_{1},\dots,\widehat{\mathbb{X}}^{(0)}_{p+q})$, where the projection matrix is $P_{Q^{(0)}}\equiv{Q^{(0)}}({Q^{(0)}}^{\intercal}Q^{(0)})^{-1}{Q^{(0)}}^{\intercal}$. In the SAR model, instrumental variables could be
$Q^{(0)}=(\mathbf{X}^{(0)},W_{1}^{(0)}\mathbf{X}^{(0)},\dots,W_{p}^{(0)}\mathbf{X}^{(0)})$
with $d=q(1+p)$ columns. We note that if the first column of $\mathbf{X}^{(0)}$ is the intercept term 1, we need to drop it in $W_{j}^{(0)}\mathbf{X}^{(0)}(j\in[p])$ to avoid multicollinearity. See Kelejian and Prucha, (1998) and Lee, (2003) for several alternative choices of instruments. The corresponding loss function for the second stage estimation is
$\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\theta\right)=\|Y^{(0)}-\widehat{\mathbb{X}}^{(0)}\theta\|_{2}^{2}.$ (8)
@@
$\widehat{\theta}_{\mathcal{A}\textit{-TranSAR}}=\underset{\theta\in\mathbb{R}^{p+q}}{\arg\min}\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\theta\right)+\lambda_{\delta}\|\theta-\widehat{\omega}\|_{1}.$ (9)
@@
$\widehat{\pi}_{j}=\underset{\pi_{j}\in\mathbb{R}^{d}}{\arg\min}\left\{\frac{1}{2n_{0}}\|{\mathbb{X}}^{(0)}_{j}-Q^{(0)}\pi_{j}\|_{2}^{2}+p_{\lambda_{\pi_{j}}}(\pi_{j})\right\},$ (10)
@@ 3.2 A transferable source detection algorithm
An intuitive approach might involve seeking a metric of similarity between each source and the target data. We compute the initial loss based on the target dataset. Subsequently, we combine the target dataset with each source dataset individually to estimate and compute the corresponding losses. If a source dataset is similar to the target dataset, we anticipate two loss values to be comparably close. Hence, we select a critical value to identify which source datasets are informative. For spatial datasets characterized by a network structure, data splitting used in Tian and Feng, (2023) can not be directly applied since it could disrupt this network integrity. As a result, we introduce the spatial residual bootstrap approach as an alternative. This idea is illustrated in details as follows.
@@
$Y^{(0,r)}=$ $\mathbb{X}^{(0)}\widehat{\theta}_{ini}+\widehat{e}^{(r)},~{}~{}\mathbf{X}^{(0,r)}=$ $\mathbf{X}^{(0)},$ (11)
for $r=1,2,3$. Let $Z^{(0,r)}=(Y^{(0,r)},\mathbf{X}^{(0,r)})$ for $r=1,2,3$. Let $Z^{(0,-r)}$ be the concatenated bootstrap samples without the $r$-th sample, for example, $Z^{(0,-1)}={({Z^{(0,2)}}^{\intercal},{Z^{(0,3)}}^{\intercal})}^{\intercal}$. Define $\mathbf{X}^{(0,-r)},\mathbb{X}^{(0,r)}$ and $\mathbb{X}^{(0,-r)}$ similarly.
Second, for each $r=1,2,3$, we fit the model in (1) on the connected bootstrap samples $Z^{(0,-r)}$ and use the resulting estimator $\widehat{\theta}^{(0,r)}$ to compute the loss value on the remaining bootstrap sample $Z^{(0,r)}$, that is
$\mathcal{L}^{[r]}(\widehat{\theta}^{(0,r)})\equiv\frac{1}{2n_{0}}\|Y^{(0,r)}-\mathbb{X}^{(0,r)}\widehat{\theta}^{(0,r)}\|_{2}^{2}.$ (12)
@@
$\hat{\sigma}=\sqrt{\frac{1}{2}\sum_{r=1}^{3}{(\mathcal{L}^{[r]}(\hat{\theta}^{(0,r)})-\mathcal{L}^{(0)})}^{2}},$ (13)
which measures the dispersion of bootstrapped loss values $\mathcal{L}^{[r]}(\widehat{\theta}^{(0,r)})$ relative to their average $\mathcal{L}^{(0)}$. Next, for each $r=1,2,3$, we combine $Z^{(0,-r)}$ with each $k$-th source dataset, fit the model in (1) on the newly combined data and compute the loss denoted by $\mathcal{L}^{(k,r)}$ on $Z^{(0,r)}$. Similarly, the average loss for the $k$-th source dataset over three-fold bootstrap samples is $\mathcal{L}^{(k)}$. Consequently, the difference between $\mathcal{L}^{(k)}$ and $\mathcal{L}^{(0)}$ provides a metric of the similarity between the $k$-th source dataset and the target data. Source datasets whose corresponding loss differences below certain threshold are recognized as transferable informative data. Following the idea of Tian and Feng, (2023), we choose $\hat{\sigma}\vee 0.01$ as the detection threshold. Then, the estimated transferable source index set $\widehat{\mathcal{A}}$ is defined as
$\widehat{\mathcal{A}}=\left\{k=1,\ldots,K\mid\mathcal{L}^{(k)}-\mathcal{L}^{(0)}\leq\left(\widehat{\sigma}\vee 0.01\right)\right\}.$ (14)
@@ 4 Residuals: $\hat{e}=Y^{(0)}-\mathbb{X}^{(0)}\hat{\theta}_{ini}$ ;
21The estimated transferable source index set: $\hat{\mathcal{A}}\leftarrow\left\{k=1,\ldots,K\mid\mathcal{L}^{(k)}-\mathcal{L}^{(0)}\leq\left(\hat{\sigma}\vee 0.01\right)\right\}$;